# Demystifying Workload Imbalances in Large Transformer Model Training over Variable-length Sequences

### v1

@@
Abstract
To optimize large Transformer model training, efficient parallel computing and advanced data management are essential. However, current methods often assume a stable and uniform training workload, neglecting imbalances in data sampling and packing that can impede performance. Specifically, data sampling imbalance arises from uneven sequence length distribution of the training data, while data packing imbalance stems from the discrepancy between the linear memory complexity and quadratic time complexity of the attention mechanism. To address these imbalance issues, we develop Hydraulis, which jointly optimizes the parallel strategies and data assignment. For one thing, we introduce large model training with dynamic heterogeneous parallel strategies in response to the sequence length variations within and across training iterations. For another, we devise a two-stage data assignment approach, which strikes a good balance in terms of the training workloads both within and across model replicas. Empirical results demonstrate that Hydraulis outperforms existing systems by 1.32-2.66$\times$.
1 Introduction
@@
In recent years, large Transformer models [80] have become pivotal in fields such as natural language processing [21, 9, 70], computer vision [10, 23, 56], video processing [7, 2, 26], and speech recognition [22, 4, 30]. Their success relies on pre-training models with vast parameters on extensive datasets [43, 25]. However, this process is both time-consuming and costly, making the acceleration of large model training a pressing concern for both academia and industry.
To move towards Artificial General Intelligence (AGI), input data are condensed into sequences of tokens to achieve a unified representation [60, 23], and models are growing larger in size [43]. Thus, large model training takes as input variable-length sequences and has substantial memory consumption. To address this, approaches from two perspectives have been developed. (a) From a parallel computing perspective, methods like data parallel [52, 87, 88, 5], tensor parallel [77, 62, 46], pipeline parallel [35, 44, 61, 49, 69, 33], sequence and context parallel [36, 54, 51, 8, 29, 27] have been proposed, each with trade-offs in latency and memory usage, and are often combined for better performance. (b) From a data management perspective, data packing techniques are widely used in the training of both language [47, 66, 6, 83, 48] and vision [20, 15] models to minimize redundant computations by combining sequences for simultaneous processing.
@@
Despite these advancements, existing training systems typically leverage these two perspectives in a decoupled manner. (a) In parallel computing, most systems overlook the sequence length variation and adopt a static, homogeneous parallel strategy [77, 72, 91, 38], which deploys all training pipelines with the same parallel scheme throughout the training. [^1] In this work, we use the term “training pipeline” (or “pipeline”) to denote a model replica that can carry out forward and backward propagation individually (i.e., a data parallel group). We use “parallel scheme” to denote the parallel configuration of a pipeline (each pipeline does not necessarily employ any parallel methods, i.e., it can be deployed on only one GPU). Given the available GPUs, a “parallel strategy” describes the collection of parallel schemes of all pipelines deployed. (b) In data management, existing works usually specify a maximum sequence length (a.k.a. context length), pack the original sequences into multiple packed ones, and distribute the packed sequences evenly across machines [66, 48].
@@
In short, current works enforce a static, homogeneous parallel strategy and a fixed-length packing on variable-length data. Nevertheless, as investigated in §3, such a decoupled design fails to take account of two key imbalances.
@@
Firstly, sequences in the real world vary in length by nature, following a long-tail distribution in general [19]. This leads to the data sampling imbalance, manifesting as variable sequence lengths within iterations and inconsistent maximum lengths across iterations. When employing a homogeneous parallel strategy, we must ensure the parallel scheme is memory-saving enough (e.g., with a high tensor parallel degree) to support the maximum sequence length, which compromises the efficiency of short sequences.
@@
The second is the data packing imbalance. Due to the attention mechanism in Transformer, there is a discrepancy between the quadratic time complexity and the linear memory complexity w.r.t. the sequence length [80, 18, 66]. However, existing works adopt fixed-length packing, achieving memory balance but leaving the training workloads imbalanced both within each pipeline and among different pipelines.
@@
Motivated by this, this work manages to address these imbalance issues by jointly optimizing the parallel and packing strategies, putting forward co-designs in both the parallel computing and data management perspectives. To this end, we introduce Hydraulis, an innovative system designed for efficient training of large Transformer models. Our major contributions are outlined as follows.
@@
(II) Disaggregating optimization and propagation with subgraphs. We present a novel disaggregation of the optimization and propagation phases in distributed training, which disentangles the divergent model partitioning brought by dynamic heterogeneous strategies. We further introduce the subgraph abstraction to unify the representation of sophisticated communication needs during propagation (for activations) and between optimization and propagation (for parameters and gradients). This allows for flexible combinations of parallel schemes and seamless transitions between parallel strategies.
@@
(III) Two-stage sequence assignment. Given a mini-batch of sequences and an arbitrary strategy, we devise a two-stage management of sequences to eliminate the data packing imbalance. In particular, we formulate an optimization problem that aims to determine how packed sequences can be balanced within each pipeline. Based on this, we also strike a good balance across different pipelines by carefully dispatching the original (unpacked) sequences. Besides, this process also estimates latency for each possible strategy, allowing us to evaluate and select the best one for each iteration.
@@
(IV) Data distribution-aware strategy proposal. Guided by the sequence assignment and for practical considerations, we further propose preparing the candidate strategies that are promising for the given training task in advance, in order to accelerate the strategy enumeration and selection for each iteration. It takes the overall sequence length distribution of the entire dataset into account, and proposes the candidate strategies via a dynamic programming process.
(V) Evaluation. Extensive experiments demonstrate that Hydraulis achieves 1.32-2.66$\times$ improvement over state-of-the-art training systems, while also offering better scalability.
2 Preliminaries
2.1 Data Sampling and Data Packing
@@
During the training of a Transformer model, each iteration samples a mini-batch of sequences with a pre-defined threshold on the total number of tokens, performs forward and backward propagation on the mini-batch, and updates the model. Besides, the model is typically trained with a pre-defined context length, representing the maximum number of tokens it can process in a single input [68]. Any sequence exceeding this limit is typically truncated to fit within the context length. But regardless of whether the original sequences are truncated or not, the lengths of the sampled sequences will inevitably exhibit variance.
@@
This variability introduces challenges when constructing batches, as sequences need to be aligned to a common length for efficient batch processing. Traditionally, this issue is addressed through data padding, where shorter sequences are padded with special tokens (e.g., [PAD] tokens) to match the length of the longest one [21, 80]. Although padding ensures uniformity in sequence length and is straightforward to implement, it leads to unnecessary computation on the padded tokens, results in wasted computational resources.
@@
Figure 1: Left: Illustration of padding and packing. Right: Comparison of the attention latency during LLaMA2 7B training on Nvidia A800 GPUs with tensor parallel (TP). Standard attention scales quadratically, while varlen achieves linear scaling relative to the sequence length (1K) before packing.
@@
(c) Subgraphs instantiation.
Figure 2: "O, P, G" represents optimizer states, parameters and gradients, respectively. "AG, RS, S, R" represents all-gather, reduce-scatter, send and receive, respectively. "To" represents the data type transfer operator (from 32- to 16-bit floating points), and "LNK" represents the link operator designed by us. (a) The overall workflow of the model states sharding technique combined with mixed-precision training. (b) Our proposed optimization-propagation disaggregation technique, which decouples the distributed strategies for the optimization phase and propagation phase, allowing for arbitrary heterogeneous parallel strategies during propagation. (c) The instantiation of Pull Parameter and Communicate Activation subgraphs during forward propagation. This involves inserting custom operators (e.g., send-receive, all-gather, LNK) to support arbitrary propagation strategies and optimization-propagation interactions.
As shown in Figure 1, an alternative to padding is data packing [47], which seeks to maximize computational efficiency by grouping sequences of varying lengths into a single longer sequence, thereby eliminating the need for padded tokens. To prevent cross-contamination between the packed sequences, positional embeddings and the attention mask must be adjusted to preserve the integrity of the attention mechanism [6, 83, 48]. Fortunately, modern techniques such as FlashAttention v2 have integrated support for block diagonal attention, which natively handles variable-length sequences (i.e., varlen) [18, 17]. Moreover, the varlen interface ensures that the attention complexity scales as $\sum{S_{i}}^{2}$, rather than $(\sum{S_{i}})^{2}$, where $S_{i}$ represents the length of the $i$-th original sequence being packed [66]. This can significantly reduce the computational overhead when using data packing, making it a more practical and efficient approach.
2.2 Parallel Methods, Schemes, and Strategies
@@
Data parallel. In data parallel (DP), the model is replicated across multiple devices, with each device processing a different subset of the input data [52, 87, 88, 5]. After computing the gradients locally, devices must synchronize their model updates, typically through a gradient reduction process, to ensure consistency across devices.
@@
Parallel scheme of one pipeline. In this work, we call each data parallel group one training pipeline, and the parallel scheme describes how a pipeline is deployed. We can broadly classify parallel methods for a pipeline into inter- and intra-op parallel, respectively.
@@
The Inter-op parallel is mainly implemented with pipeline parallel (PP). Particularly, PP divides the model into sequential stages, with each stage assigned to a different device. Data from a mini-batch is further split into micro-batches, which flow through the pipeline in a staged manner. (When training Transformer models, each micro-batch accounts for one packed sequence.) Techniques such as micro-batch scheduling and gradient accumulation are often employed, allowing multiple micro-batches to be processed efficiently [35, 44, 61, 49, 69, 33].
@@
Unlike inter-op parallel, there are several intra-op parallel methods for a pipeline. Tensor parallel (TP) partitions each layer across multiple GPUs and exchanges intermediate results (e.g., via all-gather and reduce-scatter operations) to complete the forward and backward passes [77, 62, 46]. Additionally, sequence parallel and context parallel have been explored to handle extremely long sequences, where a single sequence is partitioned across multiple devices to avoid out-of-memory (OOM) errors [36, 54, 51, 8, 29, 27].
@@
Parallel strategy. Building on the parallel scheme of a single pipeline, the parallel strategy refers to the collection of parallel schemes of all pipelines. However, existing works mainly leverage the same parallel schemes for all pipelines, resulting in a homogeneous parallel strategy. Such a homogeneous design fails to take account of the imbalances caused by data sampling and data packing, as discussed in §3.
2.3 Model States Sharding
@@
To reduce memory consumption during training, model states sharding techniques (a.k.a. ZeRO [71] and FSDP [89, 88]) are widely used. These methods extend standard data parallel (DP) by partitioning model states (i.e., parameters, gradients, and optimizer states) across multiple devices, rather than replicating them across all data parallel groups (i.e., pipelines).
Figure 3: The dot plot on the left shows the fluctuation of maximum sequence length, emphasizing inter-iteration imbalance, while the bar chart on the right presents the distribution of sequence lengths within an iteration (20th iteration), highlighting intra-iteration imbalance.
@@
The sharding technique is combined with mixed-precision training [59, 65], which uses both 16-bit floating point (FP16 or BF16) and 32-bit floating point (FP32) precisions to balance efficiency and numerical stability. Figure 2(a) illustrates an example with each pipeline deployed with a tensor parallel degree of 2. The 32-bit optimizer states and parameters (which consume more memory but are critical for ensuring convergence) are sharded across GPUs, while the 16-bit parameters and gradients (which are more memory-efficient and computationally faster) are maintained during propagation.
@@
As aforementioned, current works mainly adopt a homogeneous parallel strategy, so they usually align the sharding with the strategy, allowing distributed state transformations to be easily achieved via all-gather and reduce-scatter operations. However, since such transformations can overlap with computation, it is also feasible to adopt more nuanced, non-aligned parameter pulling and gradient pushing using customized communication strategies, which will be elaborated in §5.2.
3 Imbalances Analysis

### v2

# Hydraulis: Balancing Large Transformer Model Training via Co-designing Parallel Strategies and Data Assignment

@@
Abstract.
[…]
1. Introduction
@@
In recent years, large Transformer models (Vaswani et al., 2017) have become pivotal in fields such as natural language processing (Devlin et al., 2019; Brown et al., 2020; Raffel et al., 2020), computer vision (Carion et al., 2020; Dosovitskiy et al., 2021; Liu et al., 2021), video processing (Bertasius et al., 2021; Arnab et al., 2021; Fan et al., 2021), speech recognition (Dong et al., 2018; Baevski et al., 2020; Gulati et al., 2020) and intelligent agents (Zhou et al., 2024; Zhang et al., 2024b; Xu et al., 2025; Xi et al., 2025). Their success relies on pre-training models with vast parameters on extensive datasets (Kaplan et al., 2020; Dubey et al., 2024).
Figure 1. Sequence length distribution of two popular open-sourced datasets. We display the number of sequences and tokens within each length range, with both datasets exhibiting high variance.
Within this landscape, two key challenges arise. (a) To achieve better model ability, efforts have been made to scale models to larger sizes (Kaplan et al., 2020), making large model training substantially time-consuming and memory-intensive. (b) Meanwhile, to move towards Artificial General Intelligence (AGI), input data are often condensed into sequences of tokens to achieve a unified representation (Mikolov et al., 2012; Dosovitskiy et al., 2021; Du et al., 2024). However, this representation naturally exhibits length variability (de Vries, 2023). For instance, Figure 1 presents the sequence length distribution after tokenizing two popular large language models (LLMs) training datasets—CommonCrawl and GitHub—revealing that the sequence lengths are significantly diverse in the real world. Thus, large model training takes as input variable-length data.
To address this, approaches from two perspectives have been developed. (a) From a parallel computing perspective, methods like data parallel (Li et al., 2020; You et al., 2020; Zhang et al., 2024a; Bai, 2022), tensor parallel (Shoeybi et al., 2019; Narayanan et al., 2021b; Korthikanti et al., 2022), pipeline parallel (Huang et al., 2019; Kim et al., 2020; Narayanan et al., 2021a; Lamy-Poirier, 2023; Qi et al., 2023; Harlap et al., 2018; Guan et al., 2024), sequence and context parallel (Jacobs et al., 2023; Liu et al., 2023; Li et al., 2024; Brandon et al., 2023; Gu et al., 2024; Fang and Zhao, 2024; Wang et al., 2025a) have been proposed, and are often combined to train large-scale models. Data parallel operates at the granularity of training pipelines (i.e., data parallel groups containing model replicas that are capable of independent forward and backward propagation) and thus can typically scale by adding more pipelines, while other parallel methods (i.e., tensor, pipeline, sequence and context parallel) are optimized within a single training pipeline to balance pipeline latency and memory usage. (b) From a data management perspective, data packing techniques—combining shorter sequences into longer ones for simultaneous processing—are widely used in the training of both language (Krell et al., 2022; NVIDIA, 2023d; Bai et al., 2024; Wang et al., 2024b; Kundu et al., 2024) and vision (Dehghani et al., 2023; Choudhury et al., 2024) models to minimize redundant computations.
@@
Despite these advancements, existing training systems typically leverage these two perspectives in a decoupled manner. (a) In parallel computing, most systems overlook the sequence length variation and adopt a static, homogeneous parallel strategy (Shoeybi et al., 2019; Rasley et al., 2020; Zheng et al., 2022; Jia et al., 2019), which deploys all training pipelines (a.k.a. model replicas) with the same parallel scheme throughout the training (e.g., all pipelines have identical tensor parallel and pipeline parallel degree). [^1] In this work, we use “parallel scheme” to denote the parallel configuration of a pipeline (each pipeline does not necessarily employ any parallel methods, i.e., it can be deployed on only one GPU). Given the available GPUs, a “parallel strategy” describes the collection of parallel schemes of all pipelines deployed. (b) In data management, existing works usually specify a maximum sequence length (a.k.a. context length), pack the original sequences into multiple packed ones that do not exceed the maximum length (termed max-length packing), and distribute the packed sequences evenly across machines (NVIDIA, 2023d; Kundu et al., 2024).
@@
In short, current works enforce a static, homogeneous parallel strategy and a max-length packing on variable-length data. However, our empirical analysis (details shown in §3) reveals that this decoupled design overlooks two critical imbalances.
@@
Firstly, sequences in the real world substantially vary in length by nature (de Vries, 2023). This leads to the data sampling imbalance, manifesting as variable sequence lengths within iterations and inconsistent maximum lengths across iterations. When employing a homogeneous parallel strategy, we must ensure the parallel scheme is memory-saving enough (e.g., with a high context parallel or tensor parallel degree) to support the maximum sequence length, which compromises the efficiency of short sequences.
@@
The second is the data packing imbalance. Due to the attention mechanism in Transformer, there is a discrepancy between the quadratic time complexity and the linear memory complexity w.r.t. the sequence length (Vaswani et al., 2017; Dao et al., 2022; NVIDIA, 2023d). However, existing works adopt max-length packing, achieving memory balance but leaving the training workloads imbalanced within and across pipelines.
@@
Motivated by this, we manage to address these imbalance issues by jointly optimizing the parallel and packing strategies, putting forward co-designs in both the parallel computing and data management perspectives. To this end, we introduce Hydraulis, an innovative system designed for efficient training of large Transformer models. Our major contributions are outlined as follows.
@@
(II) Disaggregating optimization and propagation with subgraphs. To support training with dynamic heterogeneous parallel strategies, we present a novel disaggregation of the optimization phase (i.e., synchronizing gradients and updating the model) and propagation phase (i.e., executing forward and backward propagation in a pipelined manner) in distributed training. This approach maintains consistent model states sharding for optimization, and can transit to arbitrary heterogeneous parallel strategy for propagation. We further introduce subgraphs (i.e., subsets of the complete computation graph) to unify the representation of sophisticated communication needs during propagation (for activations) and optimization-propagation interactions (for parameters and gradients). This allows for flexible combinations of parallel schemes and seamless transitions between parallel strategies.
@@
(III) Two-stage sequence assignment. Given a mini-batch of sequences and an arbitrary strategy, we devise a two-stage approach to eliminate the data packing imbalance. In particular, we formulate an optimization problem that aims to determine how packed sequences can be balanced within each pipeline. Based on this, we also strike a good balance across different pipelines by carefully dispatching the original (unpacked) sequences. Besides, this process also estimates latency for each possible strategy, allowing us to evaluate and select the best one for each iteration.
@@
(IV) Data distribution-aware strategy proposal. Guided by the sequence assignment and for practical considerations, we further propose preparing the candidate strategies that are promising in advance, in order to accelerate the strategy enumeration and selection for each iteration. It takes the overall sequence length distribution of the entire dataset into account, and proposes the candidate strategies via a dynamic programming process.
[…]
2. Preliminaries
2.1. Data Sampling and Data Packing
@@
During the training of a Transformer model, each iteration samples a mini-batch of sequences with a pre-defined threshold on the total number of tokens, performs forward and backward propagation on the mini-batch, and updates the model.
Besides, the model is typically trained with a pre-defined context length, representing the maximum number of tokens it can process in a single input (Pawar et al., 2024). Any sequence exceeding this limit is typically truncated to fit within the context length. But regardless of whether the original sequences are truncated or not, the lengths of the sampled sequences will inevitably exhibit variance.
@@
This variability introduces challenges when constructing batches, as sequences need to be aligned to a common length for efficient batch processing. Traditionally, this issue is addressed through data padding, where shorter sequences are padded with special tokens (e.g., [PAD] tokens) to match the length of the longest one (Devlin et al., 2019; Vaswani et al., 2017). Although padding ensures uniformity in sequence length and is straightforward to implement, it leads to unnecessary computation on the padded tokens, resulting in wasted computational resources.
@@
Figure 2. Left: Illustration of padding and packing. Right: Comparison of the attention latency during LLaMA2 7B training on Nvidia A800 GPUs with tensor parallel (TP). Standard attention scales quadratically, while varlen achieves linear scaling relative to the sequence length (1K) before packing.
@@
Figure 3. “O, P, G” represents optimizer states, parameters and gradients, respectively. “AG, RS, S, R” represents all-gather, reduce-scatter, send and receive, respectively. “To” represents the data type transfer operator (from 32- to 16-bit floating points), and “LNK” represents the link operator designed by us. (a) The overall workflow of the model states sharding technique combined with mixed-precision training. (b) Our proposed optimization-propagation disaggregation technique, which decouples the distributed strategies for the optimization phase and propagation phase, allowing for arbitrary heterogeneous parallel strategies during propagation.
As shown in Figure 2, an alternative to padding is packing (Krell et al., 2022), which seeks to maximize computational efficiency by grouping sequences of varying lengths into a single longer sequence, thereby eliminating the need for padded tokens. To prevent cross-contamination between the packed sequences, positional embeddings and the attention mask must be adjusted to preserve the integrity of the attention mechanism (Bai et al., 2024; Wang et al., 2024b; Kundu et al., 2024). Fortunately, modern techniques such as FlashAttention v2 have integrated support for block diagonal attention, which natively handles variable-length sequences (i.e., varlen) (Dao et al., 2022; Dao, 2023). Moreover, the varlen interface ensures that the attention complexity scales as $\sum{S_{i}}^{2}$, rather than $(\sum{S_{i}})^{2}$, where $S_{i}$ represents the length of the $i$-th original sequence being packed (NVIDIA, 2023d). This significantly reduces the computational overhead of data packing, making it a more practical and efficient approach.
2.2. Parallel Methods, Schemes, and Strategies
@@
In data parallel (DP), the model is replicated across multiple devices, with each device processing a different subset of the input data (Li et al., 2020; You et al., 2020; Zhang et al., 2024a; Bai, 2022). After computing the gradients locally, devices must synchronize their model updates, typically through a gradient reduction process, to ensure consistency across devices.
Parallel scheme of one pipeline
@@
In this work, we call each data parallel group one training pipeline, and the parallel scheme describes how a pipeline is deployed. We can broadly classify parallel methods for a pipeline into inter- and intra-op parallel, respectively.
@@
The Inter-op parallel is mainly implemented with pipeline parallel (PP). Particularly, PP divides the model into sequential stages, with each stage assigned to a different device. Data from a mini-batch is further split into micro-batches, which flow through the pipeline in a staged manner. (When training Transformer models, each micro-batch accounts for one packed sequence.) Techniques such as micro-batch scheduling and gradient accumulation are often employed, allowing multiple micro-batches to be processed efficiently (Huang et al., 2019; Kim et al., 2020; Narayanan et al., 2021a; Lamy-Poirier, 2023; Qi et al., 2023; Harlap et al., 2018; Guan et al., 2024).
@@
Unlike inter-op parallel, there are several intra-op parallel methods for a pipeline. Tensor parallel (TP) partitions each layer across multiple GPUs and exchanges intermediate results (e.g., via all-gather and reduce-scatter operations) to complete the forward and backward passes (Shoeybi et al., 2019; Narayanan et al., 2021b; Korthikanti et al., 2022). Additionally, sequence parallel (SP) and context parallel (CP) have been explored to handle extremely long sequences, where a single sequence is partitioned across multiple devices to avoid out-of-memory (OOM) errors, and all-to-all or ring-style peer-to-peer communication would be necessary to perform the attention mechanism over the partitioned sequences (Jacobs et al., 2023; Liu et al., 2023; Li et al., 2024; Brandon et al., 2023; Gu et al., 2024; Fang and Zhao, 2024).
Parallel strategy
@@
Building on the parallel scheme of a single pipeline, the parallel strategy refers to the collection of parallel schemes of all pipelines. However, existing works mainly leverage the same parallel schemes for all pipelines, resulting in a homogeneous parallel strategy. Such a homogeneous design fails to take account of the imbalances caused by data sampling and data packing, which will be discussed in §3.
2.3. Model States Sharding
@@
To reduce memory consumption during training, model states sharding techniques (a.k.a. ZeRO (Rajbhandari et al., 2020) and FSDP (Zhao et al., 2023; Zhang et al., 2024a)) are widely used. These methods extend standard data parallel (DP) by partitioning model states (i.e., parameters, gradients, and optimizer states) across multiple devices, rather than replicating them across all data parallel groups (i.e., pipelines).
@@
The sharding technique is combined with mixed-precision training (Micikevicius et al., 2018; NVIDIA, 2023c), which uses both 16-bit floating point (FP16 or BF16) and 32-bit floating point (FP32) precisions to balance efficiency and numerical stability. Figure 3(a) illustrates an example with each pipeline deployed with a tensor parallel degree of 2.
The 32-bit optimizer states and parameters (which consume more memory but are critical for ensuring convergence) are sharded across GPUs, while the 16-bit parameters and gradients (which are more memory-efficient and computationally faster) are maintained during propagation.
@@
As aforementioned, current works mainly adopt a homogeneous parallel strategy, so they usually align the sharding with the parallel strategy, allowing distributed state transformations to be easily achieved via all-gather and reduce-scatter operations.
However, since such transformations can overlap with computation, it is also feasible to adopt more nuanced, non-aligned parameter pulling and gradient pushing using customized communication strategies, which is depicted in Figure 3(b) and will be elaborated in §5.2.
2.4. Graph-based Representations
The execution of a deep learning model can be represented as a computation graph—a directed acyclic graph (DAG) where nodes correspond to operators (e.g., matrix multiplication, attention) defined by users and edges indicate data dependencies. Modern deep learning frameworks, such as PyTorch (Li et al., 2020) and TensorFlow (Abadi et al., 2016), adopt this graph-based representation to encapsulate execution logic. Following this convention, we employ the same representation but extend it with a subgraph abstraction to enable dynamic heterogeneous strategies, which will be presented in §5.3.
2.5. Limitations of State-of-the-art Systems
State-of-the-art (SOTA) training systems, including Megatron-LM (Shoeybi et al., 2019; Narayanan et al., 2021b; Korthikanti et al., 2022) and DeepSpeed (Rasley et al., 2020; Jacobs et al., 2023), have incorporated diverse parallel methods (e.g, Megatron-LM supports DP, TP, PP and CP, while DeepSpeed offers DP and SP) and model states sharding techniques (e.g., Megatron-LM implements ZeRO-1, which shards the optimizer states and the FP32 parameters, while DeepSpeed supports ZeRO 1-3 to allow further sharding FP16/BF16 gradients and parameters). However, they are limited to using one static parallel strategy throughout the training task. A more recent work, HotSPa (Ge et al., 2024), enables the use of multiple pre-defined parallel strategies and supports switching between them within an iteration. This enables data in different sequence lengths within an iteration to utilize different parallel strategies while accumulating gradients for a global update.
Nonetheless, as will be analyzed in §3, various aspects of data imbalance necessitate support for dynamic heterogeneous strategies. However, all of these systems lack this capability, as they are constrained to homogeneous parallel strategies (i.e., all pipelines use the same parallel scheme) and require model states sharding to align with the parallel strategy. This is due to the following reasons: (a) Existing designs do not support parameter retrieval and gradient aggregation across heterogeneous pipelines, as they rely on homogeneous communication primitives (i.e., all-gather and reduce-scatter) but lack heterogeneous ones. (b) Balancing workload across heterogeneous pipelines is challenging due to the absence of fine-grained data assignment mechanisms and strategy design methods that take data distribution into account.
To address (a), Hydraulis adopts an optimization-propagation disaggregation architecture to facilitate sophisticated parameter retrieval and gradient aggregation, and introduces subgraphs to support arbitrary heterogeneous strategies (§5). To address (b), it proposes a two-stage sequence assignment method for load balancing (§6), and pre-generates distribution-aware heterogeneous parallel strategies before training (§7). Together, these innovations overcome the limitations of current training systems and improve training efficiency under data imbalance conditions (§8).
3. Imbalances Analysis